### v1

# Adaptive Bounding Box Uncertainties via Two-Step Conformal Prediction

Abstract
Quantifying a model’s predictive uncertainty is essential for safety-critical applications such as autonomous driving. We consider quantifying such uncertainty for multi-object detection. In particular, we leverage conformal prediction to obtain uncertainty intervals with guaranteed coverage for object bounding boxes. One challenge in doing so is that bounding box predictions are conditioned on the object’s class label. Thus, we develop a novel two-step conformal approach that propagates uncertainty in predicted class labels into the uncertainty intervals for the bounding boxes. This broadens the validity of our conformal coverage guarantees to include incorrectly classified objects, ensuring their usefulness when maximal safety assurances are required. Moreover, we investigate novel ensemble and quantile regression formulations to ensure the bounding box intervals are adaptive to object size, leading to a more balanced coverage across sizes. Validating our two-step approach on real-world datasets for 2D bounding box localization, we find that desired coverage levels are satisfied with actionably tight predictive uncertainty intervals.
1 Introduction
Safety-critical applications in domains such as autonomous transportation [36, 66] and mobile robotics [35] benefit greatly from accurate estimates of the model’s predictive uncertainty. Yet one obstacle to principled uncertainty quantification (UQ) for computer vision is the pervasive use of deep neural networks – which are often unamenable to traditional techniques for UQ. The framework of Conformal Prediction (CP) [62, 52, 2] enables a form of distribution-free UQ that is agnostic to the predictive model’s structure, rendering it well-suited for such ‘black-box’ models.
In this work, we propose a CP framework designed to quantify predictive uncertainties in multi-object detection tasks with multiple classes (see Figure 2). CP allows us to produce computationally cheap, post-hoc distribution-free prediction intervals, which come equipped with a coverage guarantee for the true bounding boxes of previously unseen objects (of known classes). Specifically, we provide users with the following safety assurance: “The conformal prediction interval covers the object’s true bounding box with probability $(1-\alpha)$ for any known object class", where $\alpha$ is an acceptable margin of error. Such a guarantee can, in the context of autonomous driving, help certify collision avoidance by steering clear of the outer-interval limits (see Figure 1), or in the case of robot picking, enforce cautious handling by demarcating a reliable grasping zone via the inner-interval limits.
Figure 1: Examples of our method for multiple classes on test images. True bounding boxes are in red, two-sided prediction interval regions are shaded in green. Produced uncertainty estimates come with a probabilistic coverage guarantee of the true boxes.
Employing strategies based on ensembling and quantile regression, we ensure that the obtained intervals are adaptive to object size: they may grow or shrink in individual dimensions to account for object variability and prediction difficulty. A challenge to the desired assurance is that constructed intervals rely on the model’s predicted class labels, which may be potentially erroneous. We thus introduce an additional conformal step over the class labels, shielding against misclassification and ensuring that downstream coverage is satisfied. That is, our two-step conformal pipeline remains theoretically and empirically valid regardless of the underlying object detector’s predictive performance for either class labels or box coordinates – the incurred costs are solely reflected in the obtained prediction interval sizes. In the experiments, we apply our methodology to multiple classes on several real-world 2D object detection datasets. We obtain bounding box prediction intervals that adhere to the desired guarantee, and are both adaptive and practically useful for downstream decision-making.
Figure 2: A diagram of our proposed two-step conformal approach. We compute conformal quantiles for both class labels and box coordinates on calibration data following the CP framework. These are then used on the predictions of a ‘black-box’ object detector for an unseen test sample to (1) form a conformal label set with guarantee (✓) which informs our box quantile choice; and (2) form a conformal prediction interval for the bounding box with guarantee (✓), thus producing a predictive uncertainty estimate with safety assurances.
To summarize, our core contribution is an end-to-end framework for safe bounding box uncertainties which is post-hoc, efficient, and generalizable. In that process, we introduce several original concepts such as (i) ensemble and quantile CP adaptations for object detection, (ii) leveraging strong class-conditional guarantees for multi-class settings, and (iii) proposing a sequential two-step approach that propagates classification uncertainties forward. Such a two-step conformal framework is, to the best of our knowledge, entirely novel for object localization.
2 Background
We begin by providing background on conformal prediction and the desired coverage guarantee, formalizing our object detection setting, and relating the conformal prediction framework to it.
2.1 Conformal prediction
Let us consider the most common setting of split CP [42], where we perform a single split to obtain separate calibration data $\mathcal{D}_{cal}=\{(X_{i},Y_{i})\}_{i=1}^{n}\sim P_{XY}$, as opposed to alternative partitioning schemes [61, 6]. If the general conformal procedure outlined in Algorithm 1 is followed, a coverage guarantee for an unseen test sample $(X_{n+1},Y_{n+1})\sim P_{XY}$ is provided in terms of a prediction set $\hat{C}(X_{n+1})$, where a finite-sample, distribution-free guarantee is given over the event of $\hat{C}(X_{n+1})$ containing $Y_{n+1}$.
Assuming the samples $\mathcal{D}_{cal}\cup\{(X_{n+1},Y_{n+1})\}$ are exchangeable – a relaxation of the i.i.d. assumption – we obtain a probabilistic guarantee that
$\mathbb{P}(Y_{n+1}\in\hat{C}(X_{n+1}))\geq 1-\alpha$ (1)
for some tolerated miscoverage rate $\alpha\in(0,1)$ [52]. The provided guarantee is marginally valid, since it holds on average across any sample $(X_{n+1},Y_{n+1})$ and set $\mathcal{D}_{cal}$ for some fixed distribution $P_{XY}$ over $\mathcal{X}\times\mathcal{Y}$. This is in contrast to the ideal scenario of conditionally valid coverage per input sample $X_{n+1}$, which has been shown to be impossible to achieve in a distribution-free manner [60, 18]. However, recent work on in-between notions of conditionality such as group- [47, 22] and feature-conditional [51] strive towards more granular guarantees.
In particular, class-conditional validity can be achieved by applying CP separately to samples from each class [10, 53, 62, 50], yielding the following guarantee:
$\mathbb{P}(Y_{n+1}\in\hat{C}(X_{n+1})|Y_{n+1}=y)\geq 1-\alpha\quad\forall y\in\mathcal{Y},$ (2)
where $\mathcal{Y}=\{1,\dots,K\}$ are distinct class labels. The class-conditional guarantee in Equation 2 is stronger and implies Equation 1, in that we aim to control some miscoverage level $\alpha$ for samples within each class. In addition, it permits setting individual miscoverage levels $\{\alpha_{y}\}_{y\in\mathcal{Y}}$ per class, if desired, and is robust to changes and imbalances in class proportions [46, 59]. Such a class-conditional guarantee is precisely what we aim to provide.
Classification and regression. Applying CP to a classification task yields conformal label prediction sets $\hat{C}_{L}(X_{n+1})\subseteq\{1,\dots,K\}$ as finite subsets of the $K$ class labels, at a target miscoverage level $\alpha_{L}$. For regression, the sets $\hat{C}_{B}(X_{n+1})\subseteq\mathbb{R}$ take the form of prediction intervals (PIs) on the target domain, at target miscoverage level $\alpha_{B}$.
2.2 Object detection
We next formalize our multi-object detection setting. Consider an input image $X\in\mathbb{R}^{H\times W\times D}$, where $H$, $W$ and $D$ correspond to image height, width and depth. For each image in $D_{cal}$ we also receive a set of tuples $(c^{1},c^{2},c^{3},c^{4},l)$, where $(c^{1},c^{2},c^{3},c^{4})\in\mathbb{R}^{4}$ are the coordinates indicating an object’s bounding box location in the image, and $l\in\{1,\dots,K\}$ represents the object’s class label.
Each tuple parameterizes an object, with a total of $O(X)$ true objects located in the image. For image $X$ we thus have responses $\{(c^{1},c^{2},c^{3},c^{4},l)_{j}\}_{j=1}^{O(X)}$. Note that the model predicts $\hat{O}(X)$ objects, and it is possible that $O(X)\neq\hat{O}(X)$. We model every object as an individual sample for our CP procedures, i.e., the same input image $X$ will produce multiple calibration samples of shape $(X,(c^{1},c^{2},c^{3},c^{4},l)_{j})$, where $j=1,\dots,O(X)$ denote the object responses.
Object detection model. For our object detector $\hat{f}$, we define two separate output heads. The probabilistic classification head is defined as the map $\hat{f}_{L}:X\mapsto(\hat{\pi}_{1},\dots,\hat{\pi}_{K})$, where $\hat{\pi}_{y}$ is the model’s estimate of the true class probability $\pi_{y}$ of some object in image $X$ belonging to class $y$. The object’s class label is then $l=\arg\max_{y\in\{1,\dots,K\}}\,\hat{\pi}_{y}$. The bounding box regression head, denoted as $\hat{f}_{B}:X\mapsto(c^{1},c^{2},c^{3},c^{4})$, maps to an object’s real-valued bounding box coordinates.
2.3 Conformal prediction for object detection
Given our multi-object detection setting, we consider a class-conditional CP approach to be particularly meaningful. It is sensible to only leverage information on detected objects of the same class, e.g., class ‘car’, to construct PIs for new objects of that class. In contrast, a general marginal approach will unintuitively also use information from unrelated classes, such as ‘person’ or ‘bicycle’.
We apply CP to the bounding boxes on a per-coordinate basis, previously denoted $(c^{1},c^{2},c^{3},c^{4})$. However, from now on let us consider the generalization to an arbitrary amount of coordinates $c^{k},\,k=1,\dots,m$. If we consider the class label $l$ within each group of objects belonging to a common class as fixed (since all objects share the same label), the response of an individual sample $(X_{i},Y_{i})$ can be interpreted as a realization of the $m$ coordinates only, i.e., we define $Y_{i}:=(c_{i}^{1},\dots,c_{i}^{m})\in\mathbb{R}^{m}$. The desired guarantee from Equation 2 is then re-interpreted as
$\mathbb{P}(\bigcap_{k=1}^{m}(c_{n+1}^{k}\in\hat{C}_{B}^{k}(X_{n+1}))|l_{n+1}=y)\geq 1-\alpha_{B}\quad\forall y\in\mathcal{Y},$ (3)
where components are indexed accordingly for a specific coordinate dimension. For instance, $\hat{C}_{B}^{k}(X_{n+1})$ is the $k$-th coordinate’s prediction interval of an object of class $y$ (its class label $l_{n+1}$ matches $y$) located in image $X_{n+1}$. Applying CP to each coordinate separately gives rise to multiple testing issues, which we address in subsection 4.1.
Practical considerations. Naively applying class-conditional CP to the box coordinates necessitates a correct class label prediction in order to satisfy validity. That is, for Equation 3 to hold, a valid PI construction requires $\hat{l}_{n+1}=l_{n+1}$ for any considered class $y\in\mathcal{Y}$. We alleviate this practically limiting dependence on the model’s classification ability using a conformal set-based classifier in section 5 (see also Figure 2). However, we of course still rely on the model’s general detection abilities – the provided guarantees only hold for true objects that are actually detected (true positives) and do not account for missed objects (false negatives), as also noted by [14, 1].
3 Related work
Many existing approaches for quantifying uncertainty in bounding box regression leverage standard UQ techniques such as Bayesian inference [12, 24, 68] and loss attenuation [21, 28, 29], or popular approximate Bayesian methods like Monte Carlo Dropout [44, 20, 38, 71] and Deep Ensembles [15, 65, 39]. These can require substantial modifications to the model architecture or training procedure, and do not provide a guarantee or statement of assurance about provided estimation quality. See Feng et al. [16] for a recent survey. A complementary line of work investigates the calibration of object detectors [45, 26, 41], which can empirically benefit our approach by improving the underlying ‘black-box’ model.
Conformal approaches have recently gained traction for computer vision and related tasks, with applications such as image classification [49, 5], geometric pose estimation [70], or tracking and trajectory planning [56, 33, 34, 40]. Yet, the domain remains comparatively unexplored given current surveys [2, 17]. Specific attempts at principled UQ for bounding boxes include using the Probably Approximatly Correct (PAC) framework to produce guarantees by composition of PAC prediction sets at multiple modelling stages [31, 43], or providing guarantees on relevant risks leveraging p-values or risk estimates obtained from concentration inequalities [3, 7, 4]. Such works differ from our two-step conformal approach in several ways, such as (i) considering different vision tasks and using different data modalities, (ii) integrating CP into complex modelling pipelines that cease to be post-hoc and model-agnostic, or (iii) leveraging non-CP methods.
CP intervals for bounding boxes have been previously considered by Andéol et al. [14, 1], whom we consider closest prior work. A key limitation in their approach is that only a single object class is considered, thus providing the most trivial form of our guarantee in Equation 3. In particular, the misclassification risk we address in section 5 does not arise for the single-class setting, where the class label is known a priori and thus $\hat{l}_{n+1}=l_{n+1}$ trivially holds. This strongly limits the practicality of their approach for real-world settings where multiple classes interact, such as in autonomous driving. Thus, our provided guarantees are theoretically stronger and more practically relevant.
In addition, we introduce several methodological improvements with good results (see subsection 6.2), such as (i) novel ensemble and quantile CP adaptations that move beyond their proposed scoring approaches, (ii) more informative two-sided intervals, as opposed to their simpler one-sided intervals, and (iii) a multiple testing correction that effectively exploits correlation structure between box coordinates, as opposed to a more naive Bonferroni [14] or $\max$-correction [1].
4 Conformal methods for box coordinates
A key modelling decision in CP is the choice of scoring function, $s:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$, for computing nonconformity scores (see Algorithm 1). We consider three choices of scoring function and subsequent PI construction for every box coordinate $k\in\{1,\dots,m\}$. Additional implementation details can be found in Appendix 0.B.
Standard conformal (Box-Std). We firstly consider the standard approach of using regression residuals as scores [52], i.e., $s(\hat{f}_{B}(X),Y)=|\hat{c}^{k}-c^{k}|$. The resulting conformal PIs are constructed as $\hat{C}^{k}_{B}(X_{n+1})=[\hat{c}^{k}_{n+1}-\hat{q}^{k}_{B},\,\,\hat{c}^{k}_{n+1}+\hat{q}^{k}_{B}]$, where $\hat{q}^{k}_{B}$ denotes the computed conformal quantile for the $k$-th coordinate. While simple, this construction only allows for non-adaptive, fixed-width PIs. Andéol et al. [14] use this approach to construct their one-sided intervals.
Conformal ensemble (Box-Ens). In order to produce more adaptive PIs, we next consider using normalized residual scores [30] of the form $s(\hat{f}_{B}(X),Y)=|\hat{c}^{k}-c^{k}|/\hat{\sigma}(X)$, where $\hat{\sigma}$ is some choice for a heuristic uncertainty estimate (i.e., without guarantees) produced by the underlying object detector. The resulting conformal PIs are constructed as $\hat{C}^{k}_{B}(X_{n+1})=[\hat{c}^{k}_{n+1}-\hat{\sigma}(X_{n+1})\,\hat{q}^{k}_{B},\,\,\hat{c}^{k}_{n+1}+\hat{\sigma}(X_{n+1})\,\hat{q}^{k}_{B}]$. By incorporating model uncertainty, the constructed PIs can be re-scaled individually per coordinate, adapting their magnitudes at test time. We may interpret this as a way of empirically conditioning on the given test sample. We employ an ensemble of object detectors and quantify $\hat{\sigma}$ as the standard deviation of the ensemble’s box coordinate predictions [27]. A joint coordinate prediction $\hat{c}^{k}$ is obtained via confidence-weighted box fusion [54].
Conformal quantile regression (Box-CQR). As a second adaptive method, we fit the score and interval construction of Conformal Quantile Regression (CQR) [48] to our setting. Additional regression heads $\hat{Q}_{B}$ are added to the object detector alongside $\hat{f}_{B}$ and trained with a quantile loss to produce lower and upper quantile predictions $\hat{Q}_{\alpha_{B}/2}$ and $\hat{Q}_{1-\alpha_{B}/2}$ for the bounding box coordinates. [^1] Note that these predictions are not the same as conformal quantiles. Under regularity conditions, these will converge asymptotically to the true conditional quantiles [23, 11] and achieve target coverage $(1-\alpha_{B})$. Following CQR, we define the scores as $s(\hat{Q}_{B}(X),Y)=\max\{\hat{Q}_{\alpha_{B}/2}(X)-c^{k},\,c^{k}-\hat{Q}_{1-\alpha_{B}/2}(X)\}$, and construct the conformal PIs as $\hat{C}^{k}_{B}(X_{n+1})=[\hat{Q}_{\alpha_{B}/2}(X_{n+1})-\hat{q}^{k}_{B},\,\,\hat{Q}_{1-\alpha_{B}/2}(X_{n+1})+\hat{q}^{k}_{B}]$. The obtained PI ensures adaptivity through the use of quantile predictions, which will differ in their distance relative to the mean coordinate prediction $\hat{c}^{k}_{n+1}$ for each sample.
4.1 Multiple testing correction
CP can also be interpreted from a hypothesis testing perspective [62, 53]. A null hypothesis for the $k$-th test sample coordinate is formed as $H_{0}:c^{k}_{n+1}=c^{k}$ for some candidate value $c^{k}\in\mathbb{R}$. Leveraging the nonconformity scores of the calibration samples as an empirical null distribution, a p-value $p(c^{k})$ is computed as the fraction of samples which conform worse than $s(\hat{c}^{k}_{n+1},c^{k})$. The value $c^{k}$ is included in the prediction set — i.e., we cannot reject $H_{0}$ — if $p(c^{k})>\alpha_{B}$. Conformalizing a set of $m$ coordinates separately and thus running $m$ nonconformity tests in parallel gives rise to multiple testing issues. Without correction, the guaranteed coverage reduces to $(1-m\,\alpha_{B})$ (see subsection 0.A.3).
A naive correction can be achieved using Bonferroni [64], since choosing $\alpha_{B}^{\prime}=\alpha_{B}/m$ will satisfy target coverage. However, the Bonferroni correction is overly conservative under positive dependency of the individual hypothesis [63], which is reasonable to assume given that the coordinates parametrize an object’s bounding box jointly. In fact, Bates et al. [8] assert that a set of conformal p-values exhibits positive dependency structure a priori since they are jointly positive regression dependent on a subset [9].
We leverage an alternate procedure by Timans et al. [58], which exploits correlation structure among box coordinates for a less conservative correction. Their suggested algorithm $\max$-rank is an adaptation of the Westfall & Young [67] permutation correction suitable for the conformal setting (see subsection 0.A.4), permitting its integration into our approach. While the use of a simple $\max$-correction has been considered for CP [1, 10], $\max$-rank operates in the scale-invariant rank space and is thus robust. In addition, it requires less computational overhead than proposed copula-based corrections [57, 37].
5 Class label prediction sets
Given our multi-class setting, the object detector is at risk of incorrectly predicting an object’s class label. This raises issues for a direct application of class-conditional CP to the object’s bounding box at test time, since we require correctly selecting the conformal box coordinate quantiles $\hat{q}^{k}_{B},\,k\in\{1,\dots,m\}$ to construct PIs that satisfy Equation 3. This limits any provided nominal safety assurances to correctly classified objects only, i.e., we require $\hat{l}_{n+1}=l_{n+1}$ for any class $y\in\mathcal{Y}$.
In order to relax the above condition on the model’s classification abilities, we introduce an additional conformal step into our modelling pipeline which preceeds the conformal bounding box construction (see Figure 2). Specifically, we consider conformalizing the model’s classifier head $\hat{f}_{L}$ to produce class label prediction sets $\hat{C}_{L}(X_{n+1})$ which guarantee correct label containment. These are subsequently used to select our box coordinate quantiles, ensuring the validity of the provided guarantee in Equation 3 is broadened to also include any incorrectly classified objects.
We achieve this using a class-conditional CP approach on the class labels with a strict label coverage guarantee of 99% (i.e., setting $\alpha_{L}=0.01$), thus approximating the condition $\hat{l}_{n+1}=l_{n+1}$ by effectively ensuring $\hat{C}_{L}(X_{n+1})\ni l_{n+1}$. The resulting two-step sequential conformal approach maintains validity regardless of the object detector’s classification or bounding box regression performance – the incurred costs are reflected in the obtained prediction set and interval sizes only. Our experiments in subsection 6.3 highlight that even under full safety assurances, our approach provides practically actionable results. We follow with a description of the employed conformal label set method and considered baselines.
5.1 Conformal class thresholding (ClassThr)
We propose using a class-conditional variant of the prediction set classifier introduced by Sadinle et al. [50], based on a similar split conformal procedure as before (see Algorithm 2). Given our probabilistic classifier $\hat{f}_{L}$, we define the scoring function $s(\hat{f}_{L}(X),y)=1-\hat{\pi}_{y}(X)$ for every class $y\in\mathcal{Y}$ and obtain per-class conformal label quantiles $\hat{q}_{L}^{y}$. The conformal label prediction set for a new object to classify is then given by $\hat{C}_{L}(X_{n+1})=\{y\in\mathcal{Y}:\hat{\pi}_{y}(X_{n+1})\geq 1-\hat{q}_{L}^{y}\}$ [^2] We invert the scores used by Sadinle et al. [50] to better reflect the notion of nonconformity, while being equivalent in procedure.. Importantly, class-conditional validity is ensured by comparing each class probability against its class-specific threshold on set inclusion.
The class-conditional guarantee stated in Equation 2 can now be provided equivalently for the object detector’s classifier head, i.e., we have that
$\mathbb{P}(l_{n+1}\in\hat{C}_{L}(X_{n+1})|l_{n+1}=y)\geq 1-\alpha_{L}\quad\forall y\in\mathcal{Y}.$ (4)
Impact on bounding box coverage guarantee. We enforce a class-conditional guarantee on the label prediction sets, since nominally requiring only a weaker marginal guarantee risks invalidating the subsequent class-conditional box coverage guarantee. If a class is systematically undercovered, we may fail to retrieve the correct box quantiles for some of its objects. Unless the class-specific label quantile is always dominated by some quantile associated with the labels present in the set, any undercoverage tendencies propagate down-stream.
The strength of the label guarantee impacts the box guarantee, since we perform both CP procedures in sequence. We observe that both guarantees are controlled in two distinct conformal procedures on the calibration data, and thus conditionally independent. It then follows for their sequential application that
$\begin{split}&\mathbb{P}(l_{n+1}\in\hat{C}_{L}(X_{n+1})\,\wedge\,\bigcap_{k=1}^{m}(c_{n+1}^{k}\in\hat{C}_{B}^{k}(X_{n+1}))\,|\,l_{n+1}=y)\\ &=\mathbb{P}(l_{n+1}\in\hat{C}_{L}(X_{n+1})|l_{n+1}=y)\,\cdot\,\mathbb{P}(\bigcap_{k=1}^{m}(c_{n+1}^{k}\in\hat{C}_{B}^{k}(X_{n+1}))|l_{n+1}=y)\\ &\hskip 150.00023pt\geq(1-\alpha_{L})(1-\alpha_{B})\quad\forall y\in\mathcal{Y}.\end{split}$ (5)
That is, a preceeding label coverage guarantee of $1-\alpha_{L}$ will nominally only assure subsequent box coverage of $(1-\alpha_{L})(1-\alpha_{B})$. In our experiments we are able to empirically approximate $(1-\alpha_{L})(1-\alpha_{B})\approx 1-\alpha_{B}$ by setting $\alpha_{L}=0.01$, thus alleviating the downstream coverage reduction. Equation 5 furthermore highlights that coverage trade-offs between objectives are possible according to sought-after assurances. For example, nominal box coverage of 90% can be achieved by both selecting $\alpha_{L}=0.05,\alpha_{B}=0.05$ and $\alpha_{L}=0,\alpha_{B}=0.1$.
5.2 Bounding box quantile selection
Obtained label prediction sets are utilized to select a conformal box coordinate quantile which provides a valid bounding box PI construction. A natural quantile selection strategy is $\hat{q}_{B}^{k}=\max\{\hat{q}_{B}^{k,y}\}_{y\in\hat{C}_{L}(X_{n+1})}\,\forall k\in\{1,\dots,m\}$, where $\hat{q}_{B}^{k,y}$ is the quantile of the $k$-th coordinate for class $y$. Using a $\max$-operator on the label set is a cautious but conservative approach, since each label in the set is regarded as equally likely. We thus incur an overcoverage cost in terms of wider PIs, which may perhaps be alleviated with a different strategy. However, we find even this simple approach to yield reasonably sized box intervals. A hypothesis testing motivation can be found in subsection 0.A.5.
Label set Box interval
Method Guarantee Size Guarantee Size
Top $*{}^{*}$ Single ${}^{\dagger}$ Small
Naive $*{}^{*}$ Small ${}^{\dagger}$ Small
ClassThr Medium Medium
Full Large Large
Table 1: Provided nominal coverage guarantees and expected empirical prediction set/interval sizes for the considered label prediction set methods on the basis of both correctly and incorrectly classified objects. $*{}^{*}$op and Naive provide a label guarantee ‘for free’ if $(1-\alpha_{L})$ is below the classifier’s accuracy level for each class. ${}^{\dagger}$op and Naive provide a box guarantee for correctly classified objects only. Naive may also satisfy both guarantees assuming practically unattainable perfect model calibration.
5.3 Label set baselines
We compare conformal class thresholding (ClassThr) to a series of alternative label prediction set constructions, whose nominal guarantees and empirically expected set sizes are outlined in Table 1.
Top singleton set (Top). We return label prediction sets that only consist of the highest probability class for every sample, i.e., $\hat{C}_{L}(X_{n+1})=\{y^{*}:\hat{\pi}_{y^{*}}(X_{n+1})=\max_{y\in\{1,\dots,K\}}\,\hat{\pi}_{y}(X_{n+1})\}$. The approach returns singleton sets and comes without nominal guarantees, and its empirical coverage relies fully on the classifier’s accuracy. The distinction to our initial condition $\hat{l}_{n+1}=l_{n+1}$, which we refer to as Oracle, is subtle: instead of ensuring correct quantile selection, we permit the use of potentially wrong quantiles to construct the box PIs.
Density level set (Naive). Assuming a perfectly calibrated classifier such that $\hat{\pi}_{y}(X)=\pi_{y}(X)\,\,\forall y\in\mathcal{Y}$, the optimal prediction set is provided by density level sets. That is, we collect all labels sorted by descending $\hat{\pi}_{y}$ until we reach probability mass $1-\alpha_{L}$. Under this assumption, prediction sets then also satisfy conditional coverage for any $X\in\mathcal{X}$ [50, 49]. While unattainable in practical settings where the classifier tends to be miscalibrated (i.e., $\hat{\pi}_{y}(X)\neq\pi_{y}(X)$), it can be considered a theoretically motivated extension of the Top baseline.
Full domain set (Full). We consider the label prediction set to be the full set of possible class labels for every sample, i.e., $\hat{C}_{L}(X_{n+1})=\mathcal{Y}$ and $|\hat{C}_{L}(X_{n+1})|=K$. In combination with our quantile selection strategy outlined below, this approach provides full label coverage for every sample, that is, it ensures $\alpha_{L}=0$. However, the cost of overly inflated label prediction sets is expected to propagate to subsequent bounding box PI construction.
We do not consider other popular conformal approaches for classification such as APS [49] or RAPS [5] since they aim at empirically improving conditional coverage under the requirements of a marginal guarantee – their advantages do not extend to a class-conditional setting as ours.
6 Experiments
For our experiments we rely on pre-trained object detection models from detectron2 [69], primarily based on a Faster R-CNN architecture and trained on COCO [32]. We consider datasets COCO (validation set), Cityscapes [13] and BDD100k [72], which contain 2D bounding box annotations and are split into calibration and test data respectively. We run our two-step conformal approach for a variety of classes, and focus our results on a coherent set of COCO classes to which class labels across all datasets can be mapped to: person, bicycle, motorcycle, car, bus and truck (see subsection 0.B.3).
Since an image may contain multiple objects, we require a pairing between an object’s true and predicted bounding boxes. Similarly to [14, 1] we perform Hungarian matching [25] based on an intersection-over-union (IoU) threshold of 0.5. Throughout, we set $\alpha_{L}=0.01,\alpha_{B}=0.1$ for a nominal box coverage of $(1-\alpha_{L})(1-\alpha_{B})\approx 90\%$, and employ $\max$-rank [58] for multiple testing correction. Results are collected over repeated trials of random calibration and test splits. Additional results, including different coverage level combinations $(\alpha_{L},\alpha_{B})$, can be found in Appendix 0.C.
Figure 3: Top: Empirical coverage levels marginally across all objects (All) and across objects from selected classes for the three conformal bounding box methods on the BDD100k dataset. Target coverage is achieved both marginally and for individual classes. Bottom: Coverage levels are stratified by object size (Small, Medium, Large), showing that Box-CQR and in particular Box-Ens provide a more balanced empirical coverage across sizes. However, this comes at the cost of slightly larger intervals, as seen when comparing $MPIW$. We also visualize target coverage (Refer to caption) and the marginal coverage distribution (Refer to caption). Displayed densities are results obtained over 1000 trials.
6.1 Metrics
Let us consider a test dataset $\mathcal{D}_{test}=\{(X_{j},Y_{j})\}_{j=n+1}^{n+n_{t}}$ of size $n_{t}$. We validate our approach by assessing our key desiderata using relevant CP metrics as described below, which jointly capture desired notions of ‘reliable’ uncertainty [2].
Validity. We ensure that our CP procedure satisfies nominal coverage guarantees by verifying its empirical coverage, which we define in generality as
$Cov=\frac{1}{n_{t}}\sum_{j=n+1}^{n+n_{t}}\mathbbm{1}[Y_{j}\in\hat{C}(X_{j})],$ (6)
where $\mathbbm{1}[\cdot]$ is the indicator function which takes the form $\mathbbm{1}[l_{j}\in\hat{C}_{L}(X_{j})]$ for label prediction sets and $\mathbbm{1}[\bigcap_{k=1}^{m}(c_{j}^{k}\in\hat{C}^{k}_{B}(X_{j}))]$ for bounding box PIs. Note that $Cov$ is a random quantity parametrized by an empirical coverage distribution and may deviate from nominal coverage levels based on factors such as calibration set size [60].
Adaptivity. To verify if target coverage is achieved by compensating undercoverage on some objects with overcoverage on others, similarly to [5, 49] we verify empirical coverage across different strata, namely object sizes. We follow the COCO detection challenge [^3] https://cocodataset.org/#detection-eval and stratify across three bounding box sizes by surface area: small ($Cov_{S}$, area $\leq 32^{2}$), medium ($Cov_{M}$, area $\in(32^{2},96^{2}]$) and large ($Cov_{L}$, area $>96^{2}$).
Efficiency. Obtained conformal prediction sets and intervals are desired to be as small as possible while still satisfying target coverage (i.e., remaining valid). We define mean set size for the label prediction sets and mean prediction interval width ($MPIW$) for the bounding box PIs as
$\frac{1}{n_{t}}\sum_{j=n+1}^{n+n_{t}}|\hat{C}_{L}(X_{j})|\quad\text{and}\quad\frac{1}{n_{t}m}\sum_{j=n+1}^{n+n_{t}}\sum_{k=1}^{m}|\hat{C}_{B}^{k}(X_{j})|.$ (7)
That is, mean set size denotes the average number of labels in the obtained sets, while $MPIW$ expresses the average interval width in terms of image pixels.
Predictive performance. We also follow standard practice and validate model performance using object detection-specific metrics from the COCO challenge, i.e., average precision across multiple IoU thresholds (see Appendix 0.C).
Two-sided box intervals One-sided box intervals
Method $MPIW$ $Cov$ $MPIW$ $Cov$
DeepEns [27] $12.31\pm 0.47$ $0.21\pm 0.01$ $74.15\pm 2.01$ $0.49\pm 0.01$
GaussianYOLO [12] $7.00\pm 0.14$ $0.08\pm 0.01$ $87.07\pm 4.25$ $0.35\pm 0.01$
Andéol et al. (Best) Not Applicable $87.62\pm 1.79$ $0.91\pm 0.01$
Box-Std (Ours) $55.47\pm 2.97$ $0.88\pm 0.02$ $85.42\pm 1.99$ $0.88\pm 0.02$
Table 2: We compare our simplest method Box-Std to Andéol et al.’s best results (see subsection 0.B.5) as well as deep ensembles (DeepEns) and GaussianYOLO, two popular UQ approaches. The former is valid (green) but only designed for one-sided intervals, while the latter two heavily undercover (red), highlighting the unreliability of obtained uncertainties due to missing guarantees. Results are for COCO across the set of selected classes and 100 trials ($\alpha_{B}=0.1$).
6.2 Comparison of bounding box methods
Empirical coverage levels stratified by class labels as well as object sizes for the three proposed bounding box methods are displayed in Figure 3 for the BDD100k dataset. We observe that the target coverage of 90% is satisfied per class, validating the class-conditional guarantee provided in Equation 3. The displayed magnitudes of variation are explained by the differences in available calibration samples per class (see subsection 0.A.6 and Table 3).
We further observe that the fixed-width PI construction of Box-Std may be wide enough to cover most small objects, but will fail to account for the magnitude of large ones, with significant undercoverage. In contrast, the adaptive nature of Box-CQR and in particular Box-Ens via its uncertainty-dependent scaling can account for varying magnitudes, achieving higher coverage for large objects at a slight loss in efficiency due to the resulting wider intervals. Whilst coverage across small objects reduces somewhat, it now intuitively aligns with observed prediction difficulty (see Table 6). That is, objects which are more challenging to predict exhibit a higher variation and chance of miscoverage. The improved balance in coverage across object sizes is a purely empirical benefit of our adaptive designs – the employed CP procedure only aims to guarantee target coverage per class and does not condition on object size.
Baseline comparisons. We validate results by comparing to Andéol et al. [1, 14], modifying our own two-sided interval methods to produce one-sided PIs, and evaluating efficiency via $MPIW$ and their box stretch metric (see subsection 0.B.5 for details). Table 2 shows that we achieve marginally tighter intervals even in their own, more restricted setting, while remaining valid. We additionally evaluate generated uncertainties via deep ensembles (DeepEns) [27] and GaussianYOLO [12], two popular UQ approaches for object detection (see section 3). Results confirm the unreliability of produced uncertainties due to their lack of guarantees, as seen by severe undercoverage in both settings. Finally, a comparison of the $\max$-rank correction to Bonferroni in Table 5 asserts that substantially tighter PIs can be obtained, while overcoverage tendencies are surpressed.
[…]
6.3 Results for the two-step approach
We compare our proposed two-step approach, i.e., employing any conformal box method with a preceeding conformal step for the class labels via ClassThr, against the proposed baselines for label prediction sets in Figure 4. Each label set method’s nominal guarantee and expected efficiency is displayed in Table 1.
In line with our expectations, only approaches using ClassThr or Full (i.e., the full label set) consistently achieve target coverage for both class labels and subsequent bounding box PIs across all three datasets. Notably, while Full results in overly conservative box PIs due to its set construction, ClassThr provides surprisingly efficient label sets (with a mean set size of at most four) which translate to actionably tight box intervals. Differences exist in the efficiency of the different conformal box methods, with Box-Ens performing notably better on BDD100k. Stratifying results by object (mis-)classification supports anticipated behaviour, in that wrongly classified objects tend to exhibit slightly higher uncertainty (see Figure 8).
Oracle relies on knowing the correct quantile and thus does not require generating a label set. While it provides nominal guarantees assuming correct label prediction and empirically satisfies them with high efficiency (i.e., narrow PIs), the condition severely limits its practicality. Top consistently undercovers the true label, and may tend to slightly undercover boxes as sample sizes increase. Naive is able to consistently maintain box coverage even though label coverage is violated, with surprisingly narrow PIs. However, additional experiments in subsection 0.C.1 showcase its sensitivity to model miscalibration, yielding it less robust than ClassThr. Yet, it is interesting that methods such as Top and Naive perform quite well empirically even under void nominal guarantees. Among others, this relates back to comments made on dominating label quantiles in subsection 5.1.
Discussion and method choice. We end by discussing the choices in selecting a suitable conformal two-step approach a practitioner may want to consider. In terms of label set methods, we note the following: if the model classifier records a high accuracy and there exists a way to externally validate predicted class labels, then Top may be a highly efficient choice. If the model is strongly calibrated and only empirical assurances are sufficient, then Naive may be a suitable selection. However, if maximal safety assurances with full nominal and empirical guarantees are desired, then ClassThr is the only safe choice. We highlight that improvement potential remains: obtained PIs using ClassThr tend to overcover, presumably due to our conservative quantile selection strategy. It may be indeed possible to obtain the same set of safety assurances with even higher efficiency using a different strategy. For instance, one could consider a weighted quantile construction on the basis of the classifier’s confusion matrix.
Regarding the choice of conformal box method, we observe that Box-Std will be the most efficient if the object detection task contains only similarly sized objects. However, if objects vary substantially in size, an adaptive approach such as Box-Ens or Box-CQR will be more suitable. One could also consider designing a CP approach which explicitly satisfies target coverage over other reasonable strata beyond classes, such as object sizes or shapes.
7 Conclusion
We present a novel procedure to quantify predictive uncertainty for multi-object detection. We leverage CP to generate uncertainty intervals with a per-class coverage guarantee for new samples. Our proposed two-step conformal approach provides adaptive bounding box intervals with safety assurances robust to object misclassification. We aim to extend our approach to 3D bounding boxes, object tracking and other detection tasks in future work, focussing on similar practically useful guarantees. Whilst further improvements can be made to achieve narrower intervals, our results are promisingly tight, paving the way for a safer deployment of vision-based systems in scenarios involving decision-making under uncertainty.
Acknowledgements
We thank members of the Bosch-UvA Delta Lab and anonymous reviewers for helpful discussions and feedback. This project was generously supported by the Bosch Center for Artificial Intelligence.
Appendix 0.A Mathematical and algorithmic details
0.A.1 Split conformal prediction
The general conformal procedure for split conformal prediction [2, 42, 62, 52] is provided in Algorithm 1 below.
Algorithm 1 Split conformal prediction
1:Input: data $\mathcal{D}\subset\mathcal{X}\times\mathcal{Y}$, prediction algorithm $\mathcal{A}$, miscoverage rate $\alpha\in(0,1)$.
2:Output: Prediction set $\hat{C}(X_{n+1})$ for test sample $(X_{n+1},Y_{n+1})$.
3:Procedure:
4:Split data $\mathcal{D}$ into two disjoint subsets: a proper training set $\mathcal{D}_{train}$ and calibration set $\mathcal{D}_{cal}=\{(X_{i},Y_{i})\}_{i=1}^{n}$.
5:Fit a prediction model on the proper training set: $\hat{f}(\cdot)\leftarrow\mathcal{A}(\mathcal{D}_{train})$.
6:Define a scoring function $s:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ applied to $\mathcal{D}_{cal}$, resulting in a set of nonconformity scores $S=\{s(\hat{f}(X_{i}),Y_{i})\}_{i=1}^{n}=\{s_{i}\}_{i=1}^{n}$. A particular score $s_{i}$ encodes a notion of dissimilarity (nonconformity) between the predicted value $\hat{f}(X_{i})$ and true value $Y_{i}$.
7:Compute a conformal quantile $\hat{q}$, defined as the $\lceil(n+1)(1-\alpha)/n\rceil$-th empirical quantile of $S$. Under exchangeability of $\mathcal{D}_{cal}\cup\{(X_{n+1},Y_{n+1})\}$, the conformal quantile $\hat{q}$ is a finite sample-corrected quantile ensuring target coverage $(1-\alpha)$ by its construction.
8:For a new test sample $(X_{n+1},Y_{n+1})$, a valid conformal prediction set for $X_{n+1}$ is given by $\hat{C}(X_{n+1})=\{y\in\mathcal{Y}:s(\hat{f}(X_{n+1}),y)\leq\hat{q}\}$. Validity refers to satisfying sample coverage with marginal probability $(1-\alpha)$, i.e., the guarantee in Equation 1. We refer to the relevant literature for the proof.
9:End procedure
Rather than a marginal guarantee, we desire the class-conditional guarantee of Equation 2, which is re-interpreted as Equation 4 for our object detection setting outlined in section 2. Thus, we may relate our problem setting to Algorithm 1 as follows:
  • $\bullet$
    A calibration sample $(X_{i},Y_{i})$ consists of an input image $X_{i}$ and an existing object in that image, parametrized as $(c^{1}_{i},\dots,c^{m}_{i},l_{i})$ for an arbitrary amount of coordinates $m$ following subsection 2.3. As we run Algorithm 1 on a per-class basis, the class label $l$ for samples within a class is fixed, and we thus define $Y_{i}:=(c^{1}_{i},\dots,c^{m}_{i})$ as the object’s coordinates.
  • $\bullet$
    Since Algorithm 1 is applied to samples within each class, the available calibration set $\mathcal{D}_{cal}$ is partitioned into smaller sets of samples matching a specific class label, i.e., we obtain sets
    $\mathcal{D}_{cal,y}=\{(X_{i},Y_{i})\in\mathcal{D}_{cal}:l_{i}=y\},\forall y\in\mathcal{Y}.$
  • $\bullet$
    Finally, the conformal procedure is performed for each of the $m$ coordinates in parallel, therefore necessating a multiple testing correction as motivated in subsection 4.1 and subsection 0.A.3. Thus, we define the final set of calibration samples for the $k$-th coordinate of an object belonging to class $y$ as
    $\mathcal{D}_{cal,y,k}=\{(X_{i},c^{k}_{i}):(X_{i},Y_{i})\in\mathcal{D}_{cal,y}\wedge c^{k}_{i}\in Y_{i}\}.$
In line with our notation in section 4, the required values to compute a nonconformity score for $(X_{i},c^{k}_{i})\in\mathcal{D}_{cal,y,k}$ are the true $k$-th coordinate $c^{k}_{i}$ and the model’s prediction $\hat{f}_{B}(X_{i})=\hat{c}^{k}_{i}$; the computed conformal quantile for $\mathcal{D}_{cal,y,k}$ is denoted $\hat{q}_{B}^{k}$; and the obtained prediction interval for a test sample’s $k$-th coordinate $(X_{n+1},c^{k}_{n+1})$ is denoted $\hat{C}^{k}_{B}(X_{n+1})$. The miscoverage rate becomes $\alpha_{B}$.
Finally, we note that assuming exchangeability of $\mathcal{D}_{cal}\cup\{(X_{n+1},Y_{n+1})\}$ also implies assuming exchangeability holds for any subsets of $D_{cal}$, such as the considered partitions $D_{cal,y,k}\subset D_{cal,y}\subset D_{cal}$.
0.A.2 Conformal class thresholding (ClassThr)
The procedure to produce class label predictions sets satisfying a coverage guarantee via conformal class thresholding (ClassThr) [50] is outlined in Algorithm 2. Note that we directly sketch the algorithm for its class-conditional variation which satisfies Equation 4.
Algorithm 2 Conformal class thresholding (ClassThr)
1:Input: data $\mathcal{D}\subset\mathcal{X}\times\mathcal{Y}$, prediction algorithm $\mathcal{A}$, miscoverage rate $\alpha_{L}\in(0,1)$.
2:Output: Prediction set $\hat{C}_{L}(X_{n+1})$ for test sample $(X_{n+1},Y_{n+1})$.
3:Procedure:
4:Split data $\mathcal{D}$ into two disjoint subsets: a proper training set $\mathcal{D}_{train}$ and calibration set $\mathcal{D}_{cal}=\{(X_{i},Y_{i})\}_{i=1}^{n}$.
5:Fit a probabilistic classifier on the proper training set: $\hat{f}_{L}(\cdot)\leftarrow\mathcal{A}(\mathcal{D}_{train})$.
6:Further partition $\mathcal{D}_{cal}$ into smaller sets of samples matching a specific class label, i.e., obtain sets $\mathcal{D}_{cal,y}=\{(X_{i},Y_{i})\in\mathcal{D}_{cal}:Y_{i}=y\},\forall y\in\mathcal{Y}$.
7:Define the scoring function $s:\mathcal{X}\times\mathcal{Y}\rightarrow[0,1],\,(\hat{f}_{L}(X),y)\mapsto 1-\hat{\pi}_{y}(X)$ for some object in image $X$ belonging to class $y$. That is, compute the object’s estimated true class probability $\hat{\pi}_{y}(X)$. Its complement encodes a notion of dissimilarity (nonconformity) between the predicted and true class probabilities.
8:Define the set of class quantiles $Q=\emptyset$.
9:Begin for $y\in\mathcal{Y}$:
10:Apply $s$ to $\mathcal{D}_{cal,y}$ to obtain a set of scores $S_{y}=\{s(\hat{f}(X_{i}),Y_{i})\}_{i=1}^{n}=\{s_{i}\}_{i=1}^{n}$.
11:Compute a conformal quantile $\hat{q}_{L}^{y}$, defined as the $\lceil(n+1)(1-\alpha_{L})/n\rceil$-th empirical quantile of $S_{y}$. Under exchangeability of $\mathcal{D}_{cal}\cup\{(X_{n+1},Y_{n+1})\}$, the conformal quantile $\hat{q}_{L}^{y}$ is a finite sample-corrected quantile ensuring target coverage $(1-\alpha_{L})$ by its construction.
12:Add the quantile to the set: $Q=Q\cup\{\hat{q}_{L}^{y}\}$.
13:End for
14:For a new test sample $(X_{n+1},Y_{n+1})$, a valid conformal prediction set for $X_{n+1}$ is given by $\hat{C}_{L}(X_{n+1})=\{y\in\mathcal{Y}:\hat{\pi}_{y}(X_{n+1})\geq 1-\hat{q}_{L}^{y}\},$ where $\hat{q}_{L}^{y}\in Q$. Validity refers to satisfying sample coverage with probability $(1-\alpha_{L})$ per class, i.e., the guarantee in Equation 4. We refer to Sadinle et al. [50] for a proof.
15:End procedure
Since Algorithm 2 is used to obtain prediction sets for class labels, we now define a calibration sample $(X_{i},Y_{i})$ as consisting of an input image $X_{i}$ and the class label of an existing object in that image, i.e., we have $Y_{i}:=l_{i}$ following the object parametrization of subsection 2.2. Similarly, we have $\hat{C}_{L}(X_{n+1})\subseteq\mathcal{Y}:=\{1,\dots,K\}$ for a given classification task with $K$ classes.
Note that the procedure in Algorithm 2 may lead to empty sets $\hat{C}_{L}(X_{n+1})=\emptyset$ in some specific cases. We alleviate any empty sets by including the highest probability label, i.e., creating a singleton set according to our baseline Top:
$\hat{C}_{L}(X_{n+1})=\{y^{*}:\hat{\pi}_{y^{*}}(X_{n+1})=\max_{y\in\{1,\dots,K\}}\,\hat{\pi}_{y}(X_{n+1})\},$
and thus $|\hat{C}_{L}(X_{n+1})|=1$. This heuristic maintains validity – the guarantee already holds for empty sets, so enlargening any prediction sets does not invalidate it. Sadinle et al. [50] also propose an ‘Accretive Completion’ algorithm to address this issue. The approach iteratively reduces the quantile magnitudes on the basis of minimizing ambiguity, filling in the null regions such that ambiguous feature spaces are better reflected. Since error control is our only concern, we opt for the simpler top-class heuristic.
Finally, ClassThr may also be motivated from a theoretical perspective. Given a perfectly calibrated classifier such that $\hat{\pi}_{y}(X)=\pi_{y}(X)\,\,\forall y\in\mathcal{Y}$, the approach can be shown to perform optimally in terms of efficiency, i.e., it produces the smallest mean set size (Sadinle et al. [50], Thm. 1).
0.A.3 Multiple testing problem
We show that running Algorithm 1 in parallel $m$ times with a global miscoverage rate $\alpha_{B}$ leads to the need for a corrected miscoverage rate on the individual level. This can be interpreted as a multiple testing correction, utilizing the hypothesis testing perspective motivated in subsection 4.1.
Recall that for a test sample $(X_{n+1},Y_{n+1})$ we have $Y_{n+1}=(c_{n+1}^{1},\dots,c_{n+1}^{m})\in\mathbb{R}^{m}$, and for its $k$-th coordinate $c_{n+1}^{k}$ the value is included in the prediction set $\hat{C}^{k}_{B}(X_{n+1})$ if $p(c_{n+1}^{k})>\alpha_{B}$. Equivalently, we may state that $c_{n+1}^{k}\notin\hat{C}^{k}_{B}(X_{n+1})$ if $p(c_{n+1}^{k})\leq\alpha_{B}$.
Starting from Equation 3 for a fixed class $y$, we then have that
$\begin{split}\mathbb{P}(\bigcap_{k=1}^{m}(c_{n+1}^{k}&\in\hat{C}^{k}(X_{n+1})))\\ &=1-\mathbb{P}(\bigcup_{k=1}^{m}(c_{n+1}^{k}\notin\hat{C}^{k}(X_{n+1})))\\ &\geq 1-\sum_{k=1}^{m}\mathbb{P}(c_{n+1}^{k}\notin\hat{C}^{k}(X_{n+1}))\\ &=1-\sum_{k=1}^{m}\mathbb{P}(p(c_{n+1}^{k})\leq\alpha_{B})\\ &\geq 1-\sum_{k=1}^{m}\alpha_{B}=1-m\alpha_{B}.\end{split}$ (8)
We observe that target coverage cannot be guaranteed globally since $1-m\alpha_{B}\leq 1-\alpha_{B}$ for any $m\in\mathbb{N}^{+}$. The Bonferroni correction naively selects $\alpha^{\prime}_{B}=\alpha_{B}/m$ to ensure $1-m\alpha^{\prime}_{B}=1-\alpha_{B}$.
Figure 5: Exact marginal coverage distributions for our three considered datasets (COCO, Cityscapes, BDD100k) on the basis of their calibration set sizes. For each distribution we additionally mark the $1\%$ and $99\%$ quantiles, as well as the desired target coverage. Left: For the conformal class label prediction sets on the basis of a target miscoverage rate $\alpha_{L}=0.01$. Right: For the conformal box prediction intervals on the basis of a target miscoverage rate $\alpha_{B}=0.1$. Obtained empirical coverage levels across experiments fall within reasonable regions of the analytically derived coverage distributions.
0.A.4 Multiple testing correction via $\max$-rank
We leverage the $\max$-rank algorithm of Timans et al. [58] to correct for multiple testing. Their procedure ensures family-wise error rate control while retaining high statistical power in settings of positively dependent hypothesis tests. For our setting, this translates to ensuring a guarantee at global coverage level $(1-\alpha_{B})$ for the full bounding box, whilst surpressing any overcoverage behaviour, i.e., providing maximally tight quantiles for each box coordinate. Again, we harness our knowledge on the positive dependency of obtained conformal p-values [8, 9], thus providing tight prediction intervals whilst maintaining coverage. In particular, Timans et al. [58] show that $\max$-rank both theoretically and empirically improves over Bonferroni.
The suggested correction is an adaptation of the Westfall & Young [67] min-P permutation correction. Intuitively, $\max$-rank collapses multiple testing dimensions into a single hypothesis test, using a composite empirical null distribution consisting of the maximum rank statistics across testing dimensions. The computed quantile over these maximum rank statistics will then ensure global coverage by design, since it controls coverage over ‘worst-case’ ranks.
Timans et al. [58] crucially relate their procedure to the hypothesis testing perspective of conformal prediction. In such a setting, each testing dimension’s empirical null distribution corresponds to the ranks of values in the computed sets of nonconformity scores $S$, i.e., for each of our box coordinates. It is shown that $\max$-rank satisfies two key requirements for its integration into a CP procedure – it preserves both exchangeability and validity. These conditions are demonstrated for the two key operations of the procedure, which are 1) operating in the rank domain of nonconformity scores, and 2) applying the $\max$-operator over these ranks. We refer to their work for further details.
0.A.5 Bounding box quantile selection
In subsection 5.2 we argue that a straight-forward quantile selection strategy is $\hat{q}_{B}^{k}=\max\{\hat{q}_{B}^{k,y}\}_{y\in\hat{C}_{L}(X_{n+1})}$, where $\hat{q}_{B}^{k,y}$ is the conformal quantile of the $k$-th box coordinate for class $y$. Then, for the $k$-th coordinate of a test sample whose true class label is indeed $y$ – which we ensure is contained in $\hat{C}_{L}(X_{n+1})$ with probability $(1-\alpha_{L})\approx 1$ – we have that $\hat{q}_{B}^{k}\geq\hat{q}_{B}^{k,y}$ and thus $\mathbb{P}(s(\hat{c}^{k}_{n+1},c^{k}_{n+1})\leq\hat{q}_{B}^{k})\geq\mathbb{P}(s(\hat{c}^{k}_{n+1},c^{k}_{n+1})\leq\hat{q}_{B}^{k,y})$. That is, using $\hat{q}_{B}^{k}$ as a thresholding quantile ensures a higher chance of score conformity, and thus is more likely to decide on inclusion into the prediction interval, at the cost of potential overcoverage.
From a hypothesis testing perspective, rather than using the p-value $p(c^{k,y})$, we base our testing decision on inclusion using $p(c^{k})=\max\{p(c^{k,y})\}_{y\in\hat{C}_{L}(X_{n+1})}$, itself a valid but conservative p-value [63]. Similarly, we then have that $p(c^{k})\geq p(c^{k,y})$ and thus $\mathbb{P}(p(c^{k})>\alpha_{B})\geq\mathbb{P}(p(c^{k,y})>\alpha_{B})$, producing a higher chance of inclusion. The coverage difference between testing with $p(c^{k})$ and $p(c^{k,y})$ then marks the incurred cost in terms of inefficiency, i.e., overly large prediction interval width.
0.A.6 Empirical coverage distribution
As mentioned in subsection 6.1, empirical coverage $Cov$ is a random quantity parametrized by a coverage distribution. For a specific randomly sampled calibration set $\mathcal{D}_{cal}$, achieved coverage may deviate from the target coverage level $(1-\alpha)$ based on factors such as calibration set size. Specifically, the exact analytical form of the coverage distribution is given by
$\mathbb{P}(Y_{n+1}\in\hat{C}(X_{n+1})|\mathcal{D}_{cal})\sim\text{Beta($n+1-l,l$)},$ (9)
[…]
In Figure 5 we display the exact coverage distributions for our three considered datasets on the basis of their calibration set sizes for both label (left) and box (right) target coverage rates. We additionally mark the $1\%$ and $99\%$ quantiles of respective distributions. We see that obtainable empirical coverage levels are in line with our experiments, where results are aggregated over multiple trials of calibration set sampling. For example, in Figure 4, where we shade the coverage distribution between the above quantiles in grey ([Uncaptioned image]), obtained coverage levels using the Oracle method are within reasonable coverage bounds, even if on the lower end. Similarly, the magnitude of variations in coverage displayed in the violin plots of Figure 3 relates directly to respective calibration set sizes. This also holds for any additional results in Appendix 0.C.
Appendix 0.B Implementation details
0.B.1 Conformal ensemble (Box-Ens)
[…]
The conformal scores for Box-Ens are defined as $s(\hat{f}_{B}(X),Y)=|\hat{c}^{k}-c^{k}|/\hat{\sigma}(X)$, where $\hat{\sigma}$ is some choice for a heuristic uncertainty estimate. Note that ‘heuristic’ refers to obtained estimates without any nominal guarantee or assurance on quality. Following the deep ensemble approach [27], we quantify $\hat{\sigma}$ as the standard deviation over box coordinate predictions, and produce a joint prediction $\hat{c}^{k}$ via weighted box fusion [54]. We detail each computation in the following.
Consider $c^{k}_{i}$ the $k$-th true coordinate of an $i$-th calibration sample to be predicted. For an ensemble of size $T$, each ensemble member produces a coordinate prediction $\hat{c}^{k}_{i,t}$ and a sample-level confidence score $\hat{s}_{i,t}$ (i.e., a confidence for the overall object prediction). The fused coordinate prediction $\hat{c}^{k}_{i}$ and uncertainty estimate $\hat{\sigma}_{i}$ are then computed as
$\hat{c}^{k}_{i}=\frac{\sum_{t=1}^{T}\hat{s}_{i,t}\,\hat{c}^{k}_{i,t}}{\sum_{t=1}^{T}\hat{s}_{i,t}}\text{ and }\hat{\sigma}_{i}=\sqrt{\frac{1}{T}\sum_{t=1}^{T}(\hat{c}^{k}_{i,t}-\bar{c}^{k}_{i})^{2}},$ (10)
where $\bar{c}^{k}_{i,t}$ is the equivalent of $\hat{c}^{k}_{i}$ with equal weights $\hat{s}_{i,t}\,\forall t=1,\dots,T$.
0.B.2 Conformal quantile regression (Box-CQR)
We modify an object detection model to regress to estimated conditional quantiles of the bounding box coordinates alongside a standard mean prediction. This is achieved by supplementing the model’s final regression output layer with additional box prediction heads, freezing all pre-trained weights, and training the additional heads with a quantile loss function, also called pinball loss [55, 23].
The loss for some quantile estimator $\hat{Q}_{\tau}$ of the $\tau$-th quantile is given by
$\mathcal{L}(y,\hat{Q}_{\tau})=\begin{cases}\tau\,(y-\hat{Q}_{\tau}(x))&\text{if }y-\hat{Q}_{\tau}(x)>0\\ (1-\tau)\,(\hat{Q}_{\tau}(x)-y)&\text{else.}\end{cases}$ (11)
It intuitively penalizes both under- and overcoverage weighted by the target quantile $\tau$, and recovers the $L1$-loss for $\tau=0.5$. Since the box heads are architecturally independent, we can train arbitrary many quantile estimators in parallel, where we obtain an individual loss $\mathcal{L}(y,\hat{Q}_{\tau})$ for each $\tau$. The final loss for model updating is the sum of all individual quantile losses.
For CQR, we require only lower and upper quantiles $\tau_{1}$ and $\tau_{2}$. If we aim for target coverage $(1-\alpha_{B})$, a reasonable choice is $\tau_{1}=\alpha_{B}/2$ and $\tau_{2}=1-\alpha_{B}/2$, since the obtained interval $[\hat{Q}_{\alpha_{B}/2},\hat{Q}_{1-\alpha_{B}/2}]$ will asymptotically achieve target coverage [23, 11]. However, in practice we require further interval scaling via CQR to obtain valid coverage in finite samples. Note that the choices for $\tau_{1}$ and $\tau_{2}$ are a modelling decision, and can in fact be tuned to produce more efficient PIs without invalidating the conformal coverage guarantee [48]. However, we only consider the simple setting with $\tau_{1}=\alpha_{B}/2$ and $\tau_{2}=1-\alpha_{B}/2$. For a desired target coverage of 90%, these correspond to $\tau_{1}=0.05$ and $\tau_{2}=0.95$.
0.B.3 Dataset splits and class mappings
We display the distribution of objects per class for the selected set of classes in Table 3. Objects are assigned to either calibration or test data based on the assignment to either split for the respective image they belong to. We randomly split the images according to the following calibration set sizes as a fraction of total available data for each dataset: 50% for COCO, 50% for Cityscapes and 70% for BDD100k.
Object class
Dataset # Images Person Bicycle Motorcycle Car Bus Truck
COCO 5000 10777 314 367 1918 283 414
Cityscapes 5000 24713 5871 895 33658 477 577
BDD100k 70000 96929 7124 3023 701507 11977 27963
Table 3: Image and object counts per dataset for the selected set of common classes.
Object class
Dataset Object size All Person Bicycle Motorcycle Car Bus Truck
COCO Small 17.68 21.70 15.96 7.58 44.60 4.98 11.27
Medium 36.88 36.29 43.09 34.09 40.70 25.34 41.78
Large 45.44 42.01 40.96 58.33 14.69 69.68 46.95
Cityscapes Small 2.90 6.56 0.49 2.99 7.34 0 0
Medium 36.87 59.55 50.77 35.05 46.81 12.36 16.67
Large 60.24 33.89 48.74 61.96 45.85 87.64 83.33
BDD100k Small 23.40 28.87 5.87 12.51 23.13 0.66 2.73
Medium 49.36 60.99 66.59 53.60 46.85 28.68 39.46
Large 38.34 10.14 27.54 33.89 30.02 70.66 57.81
Table 4: Objects stratified by size as a fraction (in %) of total object counts (Table 3) for the selected set of common classes.
In Table 4 we additionally display object counts in % stratified by size, as measured via box surface area: small (area $\leq 32^{2}$), medium (area $\in(32^{2},96^{2}]$) and large (area $>96^{2}$). While Cityscapes contains virtually no small objects, we generally observe a mix of different object sizes for various classes.
[…]
  • $\bullet$
    for Cityscapes, we map classes ‘pedestrian’ and ‘rider’ to class ‘person’;
  • $\bullet$
    for BDD100k, we map classes ‘person’ and ‘rider’ to class ‘person’.
0.B.4 Model details and parameter settings
Our primary pretrained model from detectron2 (model name X101-FPN) consists of a Faster R-CNN backbone with feature pyramid network, region proposal head, and a fully connected final bounding box prediction head, trained for $\sim 37$ epochs on the COCO training split.
Inference parameters. We identify two key parameters that filter any proposal boxes to produce the final box predictions, which we fix as follows:
  • $\bullet$
    The score parameter removes any box proposals that receive a model confidence score below a specified threshold. We fix this value at a confidence of 0.5.
  • $\bullet$
    The non-maximum suppression parameter removes any superfluous box proposals that record an IoU overlap above the specified threshold, with exception of the highest confidence box. We fix this IoU threshold at 0.6.
Quantile head training. We freeze all pretrained model weights and only train the new box prediction heads for lower and upper quantiles $\hat{Q}_{\alpha_{B}/2}$ and $\hat{Q}_{1-\alpha_{B}/2}$ with a compounded quantile loss. We set the learning rate to 0.02 and train for $\sim 3000$ COCO iterations on the COCO training split with a batch size of 16.
0.B.5 Prior work comparison
We compare our methods to the conformal bounding box approaches presented in Andéol et al. [1] – an extension of [14] – which we consider closest prior work. As mentioned in section 3, their work proposes conformal scoring methods designed for one-sided, outer prediction interval construction. We thus modify our conformal box coordinate methods (Box-Std and Box-Ens) to produce similar one-sided intervals for comparison purposes. We do not consider modifying Box-CQR, because it is not straightforward how a one-sided version of its conformal scores should be constructed.
In addition, Andéol et al. [1] only consider a single object class, thus bypassing the intricacy of capturing uncertainty in the class label predictions, as our two-step approach does. Since our comparison relates to differences in the box interval constructions only, we evaluate both their and our modified approaches across the set of classes using the correct class quantiles only (i.e., via the Oracle).
Let us once again consider the 2D bounding box setting with explicit coordinate tuples $Y=(c^{1},c^{2},c^{3},c^{4})$. To maintain notational consistency with Andéol et al. [1], we rename the coordinates by grouping them into two pairs denoting box corners as $Y:=(x_{0},y_{0},x_{1},y_{1})$. We then compare to their following proposed conformal scoring methods:
  • $\bullet$
    AddBonf. We use signed residuals to compute scores as
    $s(\hat{f}_{B}(X),Y)=(\hat{x}_{0}-x_{0},\hat{y}_{0}-y_{0},x_{1}-\hat{x}_{1},y_{1}-\hat{y}_{1})$ (12)
    and obtain outer prediction interval coordinates as
    $\hat{C}_{B}(X_{n+1})=(\hat{x}_{0}-\hat{q}_{B},\hat{y}_{0}-\hat{q}_{B},\hat{x}_{1}+\hat{q}_{B},\hat{y}_{1}+\hat{q}_{B}),$ (13)
    where $\hat{q}_{B}$ are the respective coordinate-level conformal quantiles at coverage level $(1-\alpha^{\prime}_{B})$, and $\alpha^{\prime}_{B}=\alpha_{B}/4$ is the Bonferroni correction for target coverage $(1-\alpha_{B})$.
  • $\bullet$
    MultBonf. We use the scoring function
    $s(\hat{f}_{B}(X),Y)=(\frac{\hat{x}_{0}-x_{0}}{\hat{w}},\frac{\hat{y}_{0}-y_{0}}{\hat{h}},\frac{x_{1}-\hat{x}_{1}}{\hat{w}},\frac{y_{1}-\hat{y}_{1}}{\hat{h}}),$ (14)
    where $\hat{w}=\hat{x}_{1}-\hat{x}_{0}$ and $\hat{h}=\hat{y}_{1}-\hat{y}_{0}$ are the predicted box width and height respectively. Outer intervals are then constructed as
    $\hat{C}_{B}(X_{n+1})=(\hat{x}_{0}-\hat{w}\,\hat{q}_{B},\hat{y}_{0}-\hat{h}\,\hat{q}_{B},\hat{x}_{1}+\hat{w}\,\hat{q}_{B},\hat{y}_{1}+\hat{h}\,\hat{q}_{B}),$ (15)
    where once again $\alpha^{\prime}_{B}=\alpha_{B}/4$ is corrected via Bonferroni.
  • $\bullet$
    AddMax, MultMax. Instead of using a Bonferroni correction, Andéol et al. [1] also suggest taking a $\max$-operation over coordinate scores in Equation 12 and Equation 14 respectively, resulting in a set of ‘maximal’ scores. A conformal quantile $\hat{q}_{B}$ is then computed directly at target coverage over these scores and used for each coordinate, alleviating the need for further correction. The idea is close to the $\max$-rank approach [58], but instead operates directly in the domain of scores. This can be problematic due to an improper influence of score magnitudes across dimensions. Resulting outer interval coordinates are then constructed as in Equation 13 (AddMax) and Equation 15 (MultMax) simply by replacing the quantiles.
The above approaches are compared to our following modified methods:
  • $\bullet$
    Box-Std. We use the conformal scoring approach from Equation 12 in conjunction with the $\max$-rank multiple testing correction. This is the equivalent of a one-sided signed version of the regression rsiduals emplyed by Box-Std.
  • $\bullet$
    Box-Ens. We create a one-sided version of Box-Ens by using the conformal scores from Equation 14, but using ensemble-fused coordinate predictions and normalizing by the obtained ensemble uncertainties $\hat{\sigma}$.
  • $\bullet$
    Box-Mult. We additionally consider using the conformal scores from Equation 14 directly with a $\max$-rank correction, a combination approach that can be related to the normalized scores of Box-Ens.
Box stretch metric. We implement the proposed ‘box stretch’ evaluation metric from Andéol et al. [1] that assesses the additional box surface area incurred by CP. Formally, we denote the metric as
$Stretch=\frac{1}{n_{t}}\sum_{j=n+1}^{n+n_{t}}\sqrt{\frac{\mathcal{A}(\hat{C}_{B}(X_{j}))}{\mathcal{A}(\hat{f}_{B}(X_{j}))}},$ (16)
where $\mathcal{A}(\cdot)$ is the computed surface area of the bounding box formed by the respective input, i.e., the conformal outer interval bounds, and the predicted bounding box coordinates. Ideally, we desire $Stretch$ to be close to 1.0.
Mean prediction interval width. The $MPIW$ metric is formally defined for two-sided intervals only. In order to also allow comparison using $MPIW$ alongside $Stretch$, we transform all of the above one-sided interval constructions into two-sided intervals. We do so by considering the distance of the outer interval coordinates to the box center, i.e., we place a lower interval bound at the predicted box center coordinates. Note that we also do the same for our own initially two-sided methods to allow for a fair comparison.
Appendix 0.C Additional experimental results
0.C.1 Effect of calibration on Naive and ClassThr
Figure 6: We influence model calibration via temperature scaling [19] to determine regimes of model over- and underconfidence based on $ECE$ (top). We compare Naive and ClassThr for both regimes via label coverage and label set sizes. As miscalibration through overconfidence increases (bottom left), Naive fails to satisfy target coverage. As miscalibration through underconfidence increases (bottom right), the set sizes of Naive explode. In contrast, ClassThr remains stable in respective metrics across regimes.
We aim to better understand the strong empirical coverage results observed in Figure 4 for the Naive label prediction set baseline. Since Naive is theoretically motivated from a model calibration perspective, we empirically assess its sensitivity to changes in calibration, and compare results to ClassThr. We refer to Guo et al. [19] for more details on calibration and related concepts.
Specifically, we proxy model calibration via top-class calibration, which can be evaluated using the Expected Calibration Error (ECE) metric (see Figure 6). We then influence model calibration via temperature scaling [19] – a simple logit-scaling approach – to induce regimes of model miscalibration via over- and underconfidence. We find that our object detector (temperature $T=1.0$) is generally slightly overconfident, and optimal confidence in terms of ECE lies at $T^{*}=1.3$ (Figure 6, top). We then perform our two-step conformal approach using Box-Std, and compare ClassThr to the use of Naive for multiple severities of over- and underconfidence (Figure 6, bottom). Both methods are evaluated in regards to label coverage $(1-\alpha_{L})$ (left axis) and mean set size (right axis).
We find that as miscalibration through overconfidence increases (bottom left), Naive gradually fails to satisfy target coverage of 99 %. In contrast, ClassThr is able to continuously provide desired coverage levels, at the cost of increasing set sizes for high miscalibration. As miscalibration through underconfidence increases (bottom right), the set sizes of Naive explode, tending towards the full domain (mean set size $\sim 80$). In contrast, ClassThr provides tighter coverage levels and maintains reasonable set sizes. We conclude that while Naive provides good empirical results under relatively accurate model calibration (as seen in Figure 4) it is quite sensitive to miscalibration. In contrast, ClassThr seems to provide a more robust safety assurance even under settings of over- and underconfidence, at the risk of potentially large set sizes.
0.C.2 Prior work comparison and other results
Figure 7: Comparison of our modified one-sided approaches to conformal scoring methods proposed by [1] (see subsection 0.B.5) in terms of $MPIW$ (top) and $Stretch$ (bottom). We find that our approaches are competitive with regards to both evaluation metrics.
Figure 7 displays the results for the comparison of our CP approaches to the proposed conformal methods from Andéol et al. [1] (see subsection 0.B.5). We observe that our modified one-sided approaches provide competitive performance, as measured via efficiency metrics $MPIW$ and $Stretch$, and our best approach slightly outperforms prior work.
Table 5 displays results for the conformal box procedures from section 4 across all three datasets using the Bonferroni correction rather than $\max$-rank to account for multiple testing (see subsection 4.1). In comparison to results using $\max$-rank in Figure 3, Figure 9, and Figure 10, we observe generally inferior performance. This is expressed by a higher inefficiency measured via larger $MPIW$ and a tendency to overcover beyond the target level $(1-\alpha_{B})\approx 0.9$. The effects are particularly visible for datasets COCO and Cityscapes, due to the influence of its smaller calibration set sizes.
We further ablate results of our two-step approach using ClassThr by stratifying efficiency metrics across object (mis-)classification in Figure 8 for COCO. Results confirm expected behaviour, in that the model’s higher uncertainty for more ambiguous objects – as reflected by their misclassification – tends to align with obtained larger prediction sets.
Dataset Method $MPIW$ $Stretch$ $Cov$ $Cov_{S}$ $Cov_{M}$ $Cov_{L}$
COCO Box-Std 102.6767 2.6543 0.9400 0.9992 0.9792 0.8595
Box-Ens 90.8789 1.7573 0.9358 0.9291 0.9270 0.9418
Box-CQR 90.8073 2.3136 0.9374 0.9925 0.9604 0.8764
Cityscapes Box-Std 84.7931 1.9156 0.9279 0.9779 0.8790
Box-Ens 102.2210 1.7700 0.9233 0.9208 0.9243
Box-CQR 82.5204 1.8376 0.9240 0.9545 0.8910
BDD100k Box-Std 51.4646 1.9352 0.9134 0.9986 0.9684 0.7656
Box-Ens 63.6889 1.7657 0.9084 0.8735 0.8994 0.9297
Box-CQR 51.7637 1.8438 0.9087 0.9782 0.9449 0.8004
Table 5: Metrics comparison of proposed conformal methods for box coordinates (see section 4) across the three datasets using the Bonferroni correction for multiple testing. Values are means over trials and the selected set of classes.
0.C.3 Predictive performance
Figure 8: Prediction set sizes for objects stratified by classification on COCO.
We validate the predictive performance – as measured via average precision (AP) metrics following the COCO detection challenge [^5] https://cocodataset.org/#detection-eval – for our primary pretrained object detection model X101-FPN across datasets in Table 6. Obtained scores are similar to results reported by detectron2666https://github.com/facebookresearch/detectron2/blob/main/MODEL_ZOO.md, confirming that the model performs adequately as a base predictor on which we run our conformal procedures. Other pretrained models employed for Box-Ens perform similarly.
Predictive performance metrics
Dataset AP@IoU=.50:.05:.95 AP@IoU=.75 AP@IoU=.50 AP-small AP-med AP-large
COCO 0.4521 0.4937 0.6655 0.2184 0.2781 0.4281
Cityscapes 0.4320 0.4641 0.6637 0.0270 0.0459 0.2782
BDD100k 0.3098 0.3141 0.5256 0.0745 0.1400 0.3055
Table 6: Average precision (AP) scores following the COCO detection challenge metrics for our primarily employed pretrained object detection model X101-FPN (see subsection 0.B.4). Results are the mean over our selected set of classes. The primary metric AP@IoU=.50:.05:.95 averages AP scores for 10 different IoU thresholds in $[0.5,0.95]$ with step size $0.05$. AP-small, AP-med and AP-large compute scores stratified across object sizes (see our adaptivity metric in subsection 6.1).
Figure 9: Top: Empirical coverage levels marginally across all objects (All) and across objects from selected classes for the three conformal bounding box methods on the Cityscapes dataset. Bottom: Coverage levels stratified by object size (Small, Medium, Large) and $MPIW$. We also visualize target coverage (Refer to caption) and the marginal coverage distribution (Refer to caption). Displayed densities are results obtained over 1000 trials. For interpretation see subsection 6.2.
Figure 10: Top: Empirical coverage levels marginally across all objects (All) and across objects from selected classes for the three conformal bounding box methods on the Cityscapes dataset. Bottom: Coverage levels stratified by object size (Small, Medium, Large) and $MPIW$. We also visualize target coverage (Refer to caption) and the marginal coverage distribution (Refer to caption). Displayed densities are results obtained over 1000 trials. For interpretation see subsection 6.2.
0.C.4 Different target coverage levels
Figure 11: We perform ablations on different target combinations of box coverage $(1-\alpha_{B})$ and label coverage $(1-\alpha_{L})$ for a two-step conformal approach using Box-Std and ClassThr on COCO (left), Cityscapes (middle) and BDD100k (right). We denote at each intersection of dashed lines (Refer to caption) the desired target combination. We also display obtained efficiencies, as measured via $MPIW$ for conformal box intervals and mean set size for conformal label sets. Results are averaged across the selected set of classes and 100 trials.
We evaluate a series of different target combinations of box coverage $(1-\alpha_{B})$ and label coverage $(1-\alpha_{L})$ for a two-step conformal approach using Box-Std and ClassThr on all the datasets in Figure 11. Specifically, we consider all combinations of $(1-\alpha_{B})\in\{0.85,0.9,0.95\}$ and $(1-\alpha_{L})\in\{0.8,0.9,0.99,1.0\}$. Note that the combination $(1-\alpha_{B})=0.9,(1-\alpha_{L})=0.99$ is the primary coverage goal throughout this work (see also subsection 5.1).
We compare the obtained empirical coverage levels against desired target levels, denoted by the intersections of dashed lines in the top figure row. Target coverage levels are mostly satisfied with regards to box coverage, albeit the target is reached more clearly as calibration set sizes increase, as we move from COCO (left) to Cityscapes (middle) and BDD100k (right). In contrast, target levels for label coverage are fully satisfied, and in fact tend to overcover. As calibration set sizes increase, this overcoverage tendency interestingly decreases. See also subsection 0.A.6 for more details on calibration set size.
We further display obtained efficienciy, as measured via $MPIW$ for conformal box intervals and mean set size for conformal label sets, in the bottom figure row. Two trends are immediately apparent: as we increase the requirement on box coverage, $MPIW$ increases; similarly, as we increase the requirement on label coverage, mean set size increases. Interestingly, the gap between permitting minimal label miscoverage and no miscoverage at all, i.e., $\alpha_{L}=0.01$ and $\alpha_{L}=0$, tends to increase as the number of calibration samples rises. This observation motivates our choice to select $\alpha_{L}=0.01$ as a practical, approximate way of ensuring total label coverage whilst minimally affecting the downstream box guarantees.
0.C.5 Additional visualisations
We display additional examples of conformal bounding box intervals produced using our two-step approach (see Figure 2) for COCO in Figure 12, for Cityscapes in Figure 13 and for BDD100k in Figure 14. We use ClassThr to produce class label prediction sets with 99% guaranteed coverage and either Box-Std, Box-Ens or Box-CQR (from left to right by image column) to produce conformal bounding box intervals with approx. 90% guaranteed coverage. We depict a range of different classes which align with our set of selected classes of particular interest relevant for urban driving scenes: person, bicycle, motorcycle, car, bus and truck. Note that for some images we only draw intervals for a filtered amount of classes for clearer visualisations. We observe that despite its ability to ensure probabilistic guarantees even under object misclassification, the obtained bounding box uncertainty intervals are reasonably tight, and could effectively be used for improved decision-making in many instances.
Figure 12: Examples of conformal bounding box intervals produced by our two-step approach on COCO for a mixed set of classes. Left to right by column: using ClassThr in combination with Box-Std, Box-Ens or Box-CQR. True bounding boxes are in red, two-sided prediction interval regions are shaded in green.
Figure 13: Examples of conformal bounding box intervals produced by our two-step approach on Cityscapes for a mixed set of classes. Left to right by column: using ClassThr in combination with Box-Std, Box-Ens or Box-CQR. True bounding boxes are in red, two-sided prediction interval regions are shaded in green.
Figure 14: Examples of conformal bounding box intervals produced by our two-step approach on BDD100k for a mixed set of classes. Left to right by column: using ClassThr in combination with Box-Std, Box-Ens or Box-CQR. True bounding boxes are in red, two-sided prediction interval regions are shaded in green.